OpenStudy (anonymous):

|f| - 2/3 = 5/6 what is f?

OpenStudy (anonymous):

size of whatever f is

OpenStudy (experimentx):

+3/2, -3/2

OpenStudy (campbell_st):

|f| = 9/6 f = 9/6 or -9/6